OpenStudy (anonymous):

Total of $1100 invested in two funds. One fund earned 9% profit, the other earned 4% profit. Total profit ws $84. How much ws invested in each fund?

OpenStudy (mathstudent55):

Let x = amount invested at 9% Let the rest be invested at 4%. Since the total invested was $1100, the rest of the invested money is 1100 - x. So 1100 - x was invested at 4% The profit is the interest rate times the amount invested: For the amount at 9%, the profit was 9% * x = 0.09x For the amount at 4%, the profit was 4% * (1100 - x) = 0.04(1100 - x) = 44 - 0.04x The total profit is the sum of the profits at the two rates. 0.09x + 44 - 0.44x We also know that the total profit was $84, so we write an equation to equate these two expressions: 0.09x + 44 - 0.44x = 84 Solve for x to find the profit at 9% Once you have x, plug it into x + y = 1100 to get y.